Vocaloid:

A 1000kg gondola is operated on a cable between two towers 340m apart. When the gondola is exactly between the two towers it is 100m below their height. Calculate the tension at the midpoint.

Vocaloid:

Please do not comment unless you are making a serious attempt at a solution.

Narad:

Let the tension in the cables be T By symmetry, the tension is the same in both parts of the cable. The mass = 1000 kg When the system is in equilibrium \[\sum_{}^{} Forces =0\] Therefore, \[2Tcos \theta =Mg\] By resolving in the vertical direction where the angle betweeen the cable and the vertical is \[\theta = \tan ^{-1} (170/100)\] \[g=9.81 ms ^{-2}\] \[T=(Mg)/\cos \theta \] \[=1000*9.81/\cos \theta =1000*9.81/0.507= 19348N\]

Vocaloid:

that's what I thought too at first but the answer key marks the solution as 10k N :/

Vocaloid:

@sillybilly123 would you mind taking a look at this when you get a chance? Narad and I both get ~20k Newtons but the textbook solution is 10k N

Vocaloid:

(also this is an MCAT problem where we don't have calculators so they just expect us to use 10 as g instead of 9.81 lol)

Gdeinward:

I know this may not be of help, but considering my education level here is a site that features the exact problem: https://forums.studentdoctor.net/threads/another-physics-question.1076530/

Gdeinward:

it seems to do a great job of explaining it.

Narad:

The mistake is \[2Tcos \theta =Mg\] \[\cos \theta =0.5\] \[M=10000 N\] \[T=(Mg)/(2\cos \theta )\] \[=10000/(2*0.5)=10000 N\]

sillybilly123:

always draw the FBD, even if it's obvious. Cough, splutter